### v1

# Characterization of the western Pictor A hotspot in the hard X-rays with NuSTAR

Abstract
The origin of X-ray emission from the resolved kiloparsec-scale jets of many active galactic nuclei (AGN) remains uncertain, particularly where the X-ray emission is separate from the radio-optical synchrotron component. Possible explanations include inverse-Compton upscattering of the Cosmic Microwave Background (IC-CMB) and synchrotron emission from a second electron population, alternatives which imply very different physical conditions within the jet. Until now, X-ray studies of resolved jets have been restricted to the soft X-rays (below 8 keV), often showing a hard spectral index indicating a spectral peak at higher energies. Here we present NuSTAR and Chandra observations of the nearby powerful radio galaxy Pictor A, in which we clearly detect the western hot spot at 4’ from the host galaxy, the first detection of a kpc-scale jet above 10 keV. The NuSTAR 320 keV spectrum is best fit by a powerlaw of index . Using a maximum likelihood model, we confirm previous reports of variations in the soft X-ray flux detected by Chandra over the 2000 to 2015 period, at a 6.7 level, higher than the 3.4 significance level. This rises to in the common 3-8 keV band using the full 22-year span of Chandra and NuSTAR observations. The variability of Pictor A rules out IC-CMB as the source of the X-ray emission, and the NuSTAR observations are in keeping with a synchrotron origin for the X-ray emission from the western hotspot of Pictor A, with no sign of a cutoff to the flat spectrum.
1 Introduction
A fraction of active galactic nuclei (AGN) launch large, collimated jets of relativistic plasma (Blandford & Königl, 1979). These jets can transport energy and matter up to Megaparsec distances from the central black hole in the most extreme cases (Dabhade et al., 2020). In more powerful jets, the relativistic jet can extend into the intergalactic medium (IGM), creating a bright, terminal hot spot where the kinetic energy of the jet is injected into large, extended lobes of cooling plasma. The power and extent of AGN jets is such that they can produce Mpc-scale shocks and cavities in the IGM and the intercluster medium and are thought to have a significant impact on the star-formation rate of the host galaxy (McNamara et al., 2009; Fabian, 2012).
In addition to (possible) terminal hot spots, AGN jets also generally feature discrete emitting regions called knots. Both the knots and hot spots are primarily detected at radio and sometimes IR/optical frequencies where they emit synchrotron radiation from electrons gyrating within the jet magnetic field. For this reason, jetted AGN are more ‘radio-loud’ compared to the majority of non-jetted AGN. Sources where the jet is aligned in the plane of the sky are known as radio galaxies. Radio galaxies are often divided into two classifications based on jet morphology: Fanaroff Riley type I (FR I; Fanaroff & Riley, 1974) jets tend to be shorter, less collimated, and brightest near the central engine with the jet typically terminating in large, darkened plumes of plasma, while FRII jets are highly collimated and brightest at their terminating hot spots.
Low-spatial-resolution photometric observations of jetted AGN are usually dominated by emission from the highly relativistic base of the jet (sometimes called the ‘core’ due to appearance in e.g., radio-band imaging). The focus of this study is on the emission of the extended jet, from kpc from the central engine. At these scales, the kpc-scale jet extends into and beyond the host galaxy and can be resolved in high-resolution imaging with e.g. the Hubble Space Telescope, the Very Large Array, and the Chandra X-ray Observatory.
One of the major discoveries of Chandra was the detection of ‘anomalously’ bright X-ray emission from resolved radio jets on kpc scales (Chartas et al., 2000; Sambruna et al., 2001). In some FRI jets such as M87 (Harris et al., 2003; Sun et al., 2018) and Cen A (Kraft et al., 2002; Snios et al., 2019), the X-ray emission is roughly consistent with a single spectral component extending from radio to optical to X-rays. In other, often more powerful jets (Harris & Krawczynski, 2006, 2007; Harris et al., 2010; Marshall et al., 2010) the X-ray emission is often ‘anomalous’, meaning that the X-ray flux is too bright and hard-spectrum to be an extension of the primary radio-optical synchrotron component.
While it was early on concluded that the X-ray emission expected due to synchrotron self-Compton emission is far below the observed X-ray flux (e.g., Schwartz et al., 2000), Tavecchio et al. (2000) and Celotti et al. (2001) independently showed that relativistically boosted inverse-Compton upscattering of cosmic microwave background photons (IC-CMB) could produce sufficient X-rays, provided that the kpc-scale jet is closely aligned to the line-of-sight () and remains highly relativistic (bulk Lorentz factor ). In this case, IC-CMB emission can be Doppler boosted to levels in agreement with observation.
Though often adopted as the default mechanism for high X-ray fluxes in kpc-scale jets, there is now significant evidence in the literature suggesting that IC-CMB is, at best, only capable of explaining the emission for a subset of sources. Observations incompatible with the model include high optical/UV polarization fractions in the second spectral component (Cara et al., 2013), spatial offsets between X-ray and radio emitting regions in many sources (Sambruna et al., 2004; Jester et al., 2006; Kataoka et al., 2008; Reddy et al., 2021a), and inconsistent values of and as measured from jet-counterjet flux ratios (Kataoka et al., 2008; Clautice et al., 2016; Hardcastle et al., 2016) as well as a lack of proper motions in the kpc-scale jet (Meyer et al., 2016). The IC-CMB model also predicts a significant steady GeV -ray flux (Georganopoulos et al., 2006), at a level violated by deep gamma-ray band upper limits from the Fermi/LAT for more than a dozen sources (Meyer et al., 2015, 2017; Breiding et al., 2017, 2023). Despite these challenges, IC-CMB is still frequently adapted as the default model (e.g. Kharb et al., 2012; Gentry et al., 2015). It may, in fact, be the correct explanation for anomalous X-ray emission in sources at high redshift (), where the CMB energy density is higher (e.g. Siemiginowska et al., 2003; Hodges-Kluck et al., 2021), and rare low redshift sources which satisfy the aforementioned requirements for and (e.g. Meyer et al., 2019).
The main alternative explanation is that the second, X-ray-producing component is synchrotron radiation from a second, high-energy electron population, distinct from the radio-emitting population (Hardcastle et al., 2004; Jester et al., 2006). This would require in situ particle acceleration to accelerate the electrons within the X-ray bright knots and hot spots to the necessary TeV energies as the radiating particles have cooling timescales much smaller than the transport time down the jet.
The source of the kpc-scale X-ray emission for powerful jets is a longstanding open question with the competing IC-CMB and synchrotron models often implying orders-of-magnitude differences in jet kinetic power and magnetic field strength. Determining to what extent each mechanism contributes across the X-ray jet population is thus important also for understanding the energy content of jets and their impact.
Given the unsolved origin of the X-ray emission, it is useful to attempt to extend the spectral coverage with observations in the hard X-ray range ( keV). If the emission is synchrotron in origin, locating the turnover point where the spectrum begins to fall can place constraints on the maximum electron energy. This energy range is covered by NuSTAR, but observations of kpc-scale jets are complicated by the fact that NuSTAR lacks the angular resolution (FWHM7.5”) to easily distinguish the kpc-scale jet from the core in the case of most sources. Given these limitations, the outer jet and western hot spot of FR II radio galaxy Pictor A (; 0.69 kpc/”) is nearly the only jet with the sufficient angular separation () from the core to be a viable target for hard X-ray study and the only one bright enough to be detected by NuSTAR in less than approximately 1 Ms of exposure.
Pictor A is among the most well-studied X-ray jets in the literature, and has been the subject of a number of detailed studies over the past two decades (e.g., Wilson et al., 2001; Hardcastle et al., 2004; Hardcastle & Croston, 2005; Migliori et al., 2007; Tingay et al., 2008; Marshall et al., 2010; Hardcastle et al., 2016; Thimmappa et al., 2020). The X-ray jet was first reported by Wilson et al. (2001) and was noted to be X-ray-anomalous by Hardcastle et al. (2004). Despite its bright X-ray emission, it is not a source which we would expect a significant IC-CMB component as it has a low redshift, its (approaching) western jet is not closely inclined to our line-of-sight (), and its counterjet is clearly visible (Hardcastle & Croston, 2005). A beamed IC-CMB emission origin for the X-rays from the jet or hot spot would require magnetic field strengths significantly below equipartition values (Hardcastle et al., 2004). The deepest observation of the jet and hot spot to date in the soft X-rays combined 464 ks of Chandra exposure on the western hot spot, finding a spectrum near-equally consistent with a single flat powerlaw spectrum with or a steepening, broken powerlaw spectrum with , , and a break energy at around 2 keV (Hardcastle et al., 2016). One of the motivations for our deep NuSTAR observation was to clarify the presence and location of a possible spectral break in the broad X-ray band.
The jet of Pic A is also noteworthy for the detection of variability in the X-rays. In 2010, Marshall et al. (2010) analyzed three Chandra observations taken in 2000, 2002, and 2009 of 26, 96, and 54 ks of exposure time respectively. They found that the soft X-ray flux of two features located at 34 and 49 kpc away from the core decreased significantly ( and respectively) over the course of the observations. In the later Hardcastle et al. (2016) study (with an additional 6 Chandra observations) found the hot spot X-ray flux decreased by over a time period of approximately one month, and confirmed 3 variability in the jet. Very recently, Pictor A was included in a large archival study of Chandra jet observations where at least five regions in the jet, including the hotspot, were found to be variable (Meyer et al., 2023); we further expand on the results of that study here using the same methods.
The IC-CMB mechanism requires low-energy electrons () with cooling timescales on the order of yrs (Worrall, 2009), so the observed variability in the X-ray flux is clearly incompatible with such an origin.
In this paper, we report the results of a comprehensive X-ray spectral analysis of the western hotspot of Pictor A, including new and archival NuSTAR observations totalling 348 ks. We combine this with 278 ks of archival Chandra data from 11 different observations with the Advanced CCD Imaging Spectrometer (ACIS; Garmire et al., 2003) and modeled the combined wide X-ray band spectrum ( keV) of the hot spot for the first time. We also used the maximum likelihood method of Meyer et al. (2023) to test for variability in the NuSTAR and Chandra observations, and discuss the implications of our findings for the X-ray emission origin in Pictor A.
2 Observations and data reduction
2.1 NuSTAR observations
The NuSTAR observatory consists of two independent, co-aligned X-ray telescopes, and was launched into orbit in 2012 (Harrison et al., 2013; Madsen et al., 2014). The telescopes each have an array of grazing incidence X-ray mirrors arranged in a conical approximation of a Wolter I geometry, which are designated as Optical Module A (OMA) and Optical Module B (OMB). OMA and OMB focus incoming X-rays within the FOV onto two modules of solid state detector arrays, which are designated as Focal Plane Module A (FPMA) and Focal Plane Module B (FPMB).
NuSTAR first observed the western hot spot of Pictor A in 2015 with an exposure of 109 ks. A second observation was proposed and completed in 2021, consisting of 3 pointings totaling 302 ks of exposure. We group these later observations into a single ‘epoch’, or collection of observations conducted within a short timeframe of one another ( 1 wk). The details of the 2 NuSTAR epochs and their constituent observations which are used in this study are included in Table 1, which lists the project code, observation date, PI, the epoch grouping, and the exposure time before and after filtering.
Instrument | Epoch | Project | Date Obs. | Exposure | Filtered | PI |
Code | (YYYY-MM-DD) | (ks) | Exposure (ks) | |||
NuSTAR | 1 | 60101047002 | 2015-12-03 | 109.5 | 92.82 | Croston |
2 | 60701064002 | 2021-11-28 | 126.2 | 104.1 | Meyer | |
2 | 60701064004 | 2021-12-01 | 37.59 | 32.39 | Meyer | |
2 | 60701064006 | 2021-12-02 | 138.8 | 118.6 | Meyer | |
Total | - | - | 412.1 | 347.9 | - | |
Chandra | 1 | 346 | 2000-01-18 | 25.83 | 25.07 | Wilson |
2 | 3090 | 2002-09-17 | 46.36 | 31.23 | Wilson | |
2 | 4369 | 2002-09-22 | 49.12 | 34.53 | Wilson | |
3 | 11586 | 2009-12-12 | 14.26 | 12.21 | Hardcastle | |
3 | 12039 | 2009-12-07 | 23.74 | 21.95 | Hardcastle | |
3 | 12040 | 2009-12-09 | 17.32 | 15.62 | Hardcastle | |
4 | 15580 | 2012-11-08 | 10.48 | 9.97 | Hardcastle | |
5 | 14222 | 2014-01-17 | 45.38 | 44.03 | Hardcastle | |
6 | 14223 | 2014-04-21 | 50.14 | 42.20 | Hardcastle | |
7 | 16478 | 2015-01-09 | 26.82 | 23.75 | Hardcastle | |
7 | 17574 | 2015-01-10 | 18.62 | 17.34 | Hardcastle | |
Total | - | - | 328.1 | 277.9 | - |
We downloaded the raw NuSTAR data from HEASARC and applied the standardized procedure to reduce and image the data using the nupipeline command from the NuSTAR Data Analysis Software (NuSTARDAS) v2.1.0111https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_swguide.pdf, which is included in the HEASARC software package (HEAsoft) v6.29222https://heasarc.gsfc.nasa.gov/docs/software/heasoft/developers_guide/. In particular, we excised time intervals where background radiation and contamination from the South Atlantic Anomaly (SAA) was significant using the nupipeline settings SAA = STRICT and TENTACLE = YES. At this stage, we filtered the data to the energy range of 3-78 keV.
Once the data were filtered, we created a circular extraction region of radius 50” at the centroid of the hot spot. We used the nuproducts pipeline to generate high-level data products for the source region, including extracted photon counts spectral files (PHAs), auxiliary response files (ARFs), response matrix files (RMFs), and exposure maps. This process was repeated for both the FPMA and FPMB images from each NuSTAR observation, giving us 8 total sets of NuSTAR data.
For visual study of the jet and environs, we created a hard X-ray image using the combined data of all available NuSTAR observations. We used the nuskybgd code developed by Wik et al. (2014) to generate background-subtracted images (see Section 2.3 for more details). We used the background-subtracted images and the CIAO commands dmstat and acrosscorr to calculate the centroid position of the auto-correlated image of the earliest FPMA observation of the X-ray core, calculating the offsets in the centroid positions of later observations, and reprojecting the later observations such that the X-ray core is aligned. We used the same offsets to align the exposure maps generated by nuproducts. We co-added all our aligned NuSTAR observations and divided the resulting deep photon image by the similarly co-added exposure map to obtain a hard X-ray flux image of Pictor A. Lastly, we applied a Gaussian smoothing filter on the image at a factor of 2 pixels (4.8”) to produce our final image, which is shown in Figure 2, with Chandra contours overlaid.
The peak of the hot spot emission profile as seen by NuSTAR is slightly offset from the peak observed by Chandra, approximately 4.7” downstream along the jet axis. This offset places the hard X-ray centroid beyond even the radio peak, which is estimated to be downstream of the Chandra position by approximately 1” (Hardcastle et al., 2016). We estimate the uncertainty in the centroid position to be (dx,dy) = (1.72”,2.42”), which we take from the root mean square correction obtained from cross-correlations on the NuSTAR images with the acrosscorr command. However, this offset is most likely due to an asymmetry in the effective PSF combined with the alignment uncertainty. The deconvolved NuSTAR images (see Section C) show a peak consistent with the Chandra images.
2.2 Chandra observations
The Chandra X-ray Observatory was launched in 1999 as a part of NASA’s Great Observatories program (Weisskopf et al., 2000). The ACIS-S CCD detector (Garmire et al., 2003) onboard has an unrivaled subarcsecond angular resolution in the 0.3 - 10.0 keV energy range, making it ideal for observing kiloparsec-scale extragalactic X-ray jets. Chandra has observed Pictor A in 16 separate pointings, from the initial discovery of the large-scale X-ray jet in 2000 (Wilson et al., 2001) to subsequent observations of the jet and hot spot as recently as 2015 (Hardcastle et al., 2004; Hardcastle & Croston, 2005; Hardcastle et al., 2016). A few observations were excluded from this analysis due to extremely short exposure. Similar to our NuSTAR observations, we grouped observations conducted within week of one another into single epochs. Ultimately, we made use of use 11 Chandra observations which are grouped into 7 epochs and totaling a nominal 330 ks of exposure time. A full list of the Chandra observations used in this analysis, their epoch groupings, PI, project code, observation date, and exposure time before and after filtering can be found in Table 1, alongside the NuSTAR epochs and observations described in the previous section.
All observations were reduced using CIAO version 4.1.2 (Fruscione et al., 2006) and the calibration database CALDB v. 4.9.2. Each observation was first reprocessed using the chandra_repro and filtered to remove all photons outside of the 0.4-8 keV range to minimize background contamination. We use the lc_clean task to bin the good time interval (GTI) data with a bin time of 259.28 s and excise intervals with background flares, which we define as a photon count rate which deviates from the mean count rate by more than 2. We then used this filtered data to produce the 11 sub-pixel images centered on the core of the X-ray jet, one for each epoch, downsizing by a factor of 5 for a pixel scale of 0.0984”. After filtering, the total exposure time is approximately 278 ks.
We also generated a combined soft X-ray image using all Chandra data reduced for this study. We again used cross-correlation to precisely locate the hot spot centroid and structure to correct for astrometric offsets between observations and the exposure maps generated for each observation using the CIAO tool mkexpmap. Finally, we combined the aligned images for each Chandra epoch using the CIAO tool dmmerge and divided by the combined exposure map to obtain the final soft X-ray flux image of Pictor A. The final image is presented in Figure 3, with NuSTAR contours overlaid.
2.3 NuSTAR background
The background radiation spectrum for NuSTAR observations is complex, including multiple different components which dominate at different energies. Additionally, the background radiation is spatially nonuniform and varies significantly according to the position on the detector.
The background spectra for each module and pointing were modeled and fit using the IDL-based nuskybgd code developed by Wik et al. (2014), which maps both the spectral and spatial variations of the background emission across the entire image. The full background model is detailed in Appendix A and the best-fit normalization for each variable parameter for each NuSTAR observation is recorded in Supplemental Table 1 in the online version of this article. For all observations, /DOF had a value between 0.9 and 1.1, indicating that the background was well fit by the model. Figure 4 shows a visual representation of the complete background model, including the different individual components, taken from Wik et al. (2014) alongside the background spectra observed in one of our NuSTAR observations.
We used nuskybgd to extract and fit the background spectrum and make a map of the background emission over the detector coordinates for each NuSTAR observation. After fitting the background, we provided nuskybgd with our source region and created a model file containing the appropriate parameter values for the background spectrum at those detector coordinates.
2.4 Chandra background
For each Chandra observation, we extracted the background photon counts from a large elliptical region far from the hot spot and bright X-ray core with minor and major axes of 70” and 65” respectively. We used this to generate a background spectral file corresponding to each Chandra observation of the hot spot. Whenever fitting Chandra data in XSPEC, we use the CSTAT statistic which defaults to WSTAT when background data is loaded but not modeled in XSPEC. We bin the spectra such that there are at least 5 counts in the background spectra for each bin to avoid any problematic biases.
2.5 Spectral modeling of the western hot spot
We used XSPEC v12.10.1f and its python interface PyXspec v 2.1.1 for all X-ray spectral analysis. We chose the Poisson-based cstat statistic (Cash, 1979) for our parameter estimation statistic and the Anderson-Darling (ad) test statistic to evaluate goodness-of-fit. We chose this method to avoid the problematic biases that arise from trying to bin and fit the spectrum using the statistic (Andrae et al., 2010). We fit the spectra from Chandra and NuSTAR separately, then performed a final fit using the combined data set.
For NuSTAR observations, we use a 50” radius circular region with its center located at the centroid of the western hot spot as our source region. We extracted the source counts from this region for both the FPMA and FPMB image and filtered to an energy range of 3-20 keV. Based on the results of our analysis using nuskybgd across the 2 epochs of NuSTAR data, we estimate that approximately 3192 counts out of 3542 total counts () in the energy band of 20-78 keV are the result of background radiation. Therefore, we ignored this higher energy band in our analysis.
When modeling the source spectrum for each NuSTAR epoch, we combined FPMA and FPMB data, and introduced a constant factor to our spectral model (constant parameter). For the FPMA observations, this constant was fixed to 1, while for FPMB observations, it was free to vary between 0.9 and 1.1. This factor was included to account for the unknown normalization difference between the two detectors which typically varies between (Madsen et al., 2015). After fitting each NuSTAR epoch separately, we performed a total fit using all reduced NuSTAR data.
For Chandra observations, we extracted the source counts from an elliptical region with minor and major axes of 6” and 9” respectively and its center located at the centroid of the western hot spot. Just as we did with NuSTAR, we fit each Chandra epoch separately, and then fit the combined Chandra data in an energy range of 0.5-7 keV.
Model | XSPEC | Model Formula | Parameter | Param. | Prior Type | Prior | Prior |
Name | Model | [] | Name | No. | Min. | Max. | |
Powerlaw | powerlaw | N | 2 | Log Unif. | -10 | 0 | |
PhoIndex | 1 | Unif. | 1 | 3 | |||
Broken powerlaw | bknpower | N | 5 | Log Unif. | -10 | 0 | |
PhoIndx1 | 1 | Unif. | 1 | 3 | |||
High-energy cutoff powerlaw | cutoffpl | N | 3 | Log Unif. | -10 | 0 | |
PhoIndex | 1 | Unif. | 1 | 3 | |||
HighECutoff | 3 | Unif. | |||||
Log-parabola | logpar | N | 4 | Log Unif. | -10 | 0 | |
alpha | 1 | Unif. | 1 | 3 | |||
beta | 2 | Unif. | 0 | 1 | |||
Thermal plasma | apec | Smith et al. (2001) | N | 16 | Log Unif. | -10 | 0 |
kT | 1 | Unif. | 0 | 30 | |||
Double broken powerlaw | bkn2power | N | 6 | Log Unif. | -10 | 0 | |
PhoIndx1 | 1 | Unif. | 1 | 3 | |||
PhoIndx2 | 3 | Unif. | 0 | 4 | |||
BreakE1 | 2 | Unif. | 5 | ||||
BreakE2 | 2 | Unif. | 5 |
Finally, to maximize our photon statistics and achieve the best constraints on the X-ray spectrum over the wide X-ray band (0.5-20 keV), we fit the combined spectrum of Chandra and NuSTAR data. We reintroduce the constant factor used to account for offsets between the NuSTAR FPMA and FPMB detectors, again freezing it at a value of 1 for the FPMB data and allowing it to vary from 0.9 to 1.1 for FPMA data. We also apply a constant factor to our Chandra data to similarly account for the offsets between NuSTAR and Chandra data (Madsen et al., 2017). We allow this second constant to vary in a range of 0.1-10, given the possibility of intrinsic source variability.
To model the spectrum for a given epoch or epochs, we loaded the corresponding background model or spectral files created through the processes detailed in 2.3 and 2.4. We evaluated the following six spectral models: a powerlaw (zpowerlw), a broken powerlaw (zbknpower), a double broken powerlaw (bkn2pow), a log-parabola (zlogpar), a cutoff powerlaw (zcutoffpl), and a thermal plasma model (apec). For each model, we assumed a redshift of and applied a photoelectric absorption model (phabs) with a fixed Hydrogen column density value of nH = (HI4PI Collaboration et al., 2016). For the thermal plasma model, we additionally assumed a metallicity of 0.3 times solar values.
We used the Bayesian X-ray Analysis (BXA) software from Buchner (2016), which connects a nested sampling algorithm (UltraNest; Buchner, 2021)) with the Python version of XSPEC, pyXspec (Gordon & Arnaud, 2021). We found this to be the most effective and stable method of spectral fitting. BXA also enables us to easily compare non-nested models using the Bayes’ evidence (Z), the likelihood integrated over the parameter space (Knuth et al., 2014). The tested models reflect the possibility of X-ray emission from the hot spot could potentially arise from inverse-Compton scattering, synchrotron radiation, SSC, thermal emission, or any combination of these processes.
Inst. | Proj. Code | Src. Counts | Bkg. Counts | Src. Exposure | Bkg. Exposure | ECF |
(ph) | (ph) | ( cms) | ( cms) | |||
NuSTAR | 60101047002 | 2127 | 1304 | 64.4 | - | 0.677 |
60701064002 | 3172 | 1510 | 109 | - | 0.671 | |
60701064004 | 952 | 441 | 34.2 | - | 0.670 | |
60701064006 | 3531 | 1638 | 123 | - | 0.670 | |
Chandra | 346 | 2969 | 160 | 11.9 | 12.1 | 0.973 |
3090 | 3229 | 156 | 12.5 | 12.3 | 0.982 | |
4369 | 3573 | 359 | 13.7 | 13.5 | 0.980 | |
11586 | 1105 | 104 | 3.97 | 3.89 | 0.979 | |
12039 | 1890 | 212 | 7.13 | 7.00 | 0.982 | |
12040 | 1411 | 148 | 5.07 | 5.06 | 0.977 | |
15580 | 825 | 54 | 2.86 | 2.49 | 0.978 | |
14222 | 3090 | 256 | 11.2 | 11.9 | 0.980 | |
14223 | 2380 | 194 | 7.88 | 8.42 | 0.981 | |
16478 | 1350 | 148 | 5.25 | 5.99 | 0.984 | |
17574 | 1014 | 115 | 3.61 | 4.26 | 0.976 | |
We defined our prior distributions such that the analysis spans the entire parameter space. In particular, we defined the priors for our spectral parameters (e.g. ) and energy parameters to sample a uniform distribution across all physically reasonable values. The uninformative priors are thought to minimize bias in the posterior distribution. We report the relevant prior distributions for all the free parameters of the evaluated models in Table 2. This table includes the model name, the XSPEC model name, the model formula, the XSPEC parameter names, the type of prior for the parameter (either uniform or log-uniform), and the minimum and maximum value of the prior distribution.
2.6 Variability analysis
In previous work, Meyer et al. (2023) introduced a maximum-likelihood based statistical test to analyze a series of CCD observations of point sources for variability where the null hypothesis is a non-variable Poisson source rate and a variable background rate. The test uses a maximum likelihood model to return a p-value for a set of single-region observations under the null hypothesis. Previously, this method was used to conduct a variability survey of the large-scale, X-ray jet population using Chandra archival data of 53 sources, including Pictor A (Meyer et al., 2023). While Chandra’s high angular resolution make it an ideal partner for general tests of resolved jet variability, the test is generally applicable to e.g. other observatories with CCD detectors. Here we applied the variability test to both our Chandra and NuSTAR data on Pictor A.
The maximum likelihood test is described in full in Meyer et al. (2023). Here, we will only briefly describe the necessary elements of the test. For each observation entered into our analysis, we required the following data: (1) the total number of photon counts in the source region, (2) the number of background photons in the source region, (3) the exposure in the source region, and (4) the encircled counts fraction (ECF). For the analysis of the Chandra alone, we extracted this data for an energy range of 0.4-8 keV, and for NuSTAR alone, we extracted it for the full 3-78 keV energy range. We also separately extracted the 3-8 keV counts for a combined analysis using both observatories. We list the values for this data for all observations used in the variability analysis are given in Tables 3 and 4. Using this data, the maximum likelihood model returns an estimate of the source flux, , under the steady-rate assumption and also the probability of observing the data under the null hypothesis. A low p-value indicates that the null hypothesis can be rejected and that the intrinsic source rate is therefore likely to be variable.
Inst. | Proj. Code | Src. Counts | Bkg. Counts | Src. Exposure | Bkg. Exposure | ECF |
(ph) | (ph) | ( cms) | ( cms) | |||
NuSTAR | 60101047002 | 978 | 426 | 21.6 | - | 0.734 |
60701064002 | 1412 | 304 | 35.5 | - | 0.745 | |
60701064004 | 396 | 89 | 11.0 | - | 0.736 | |
60701064006 | 1497 | 333 | 40.3 | - | 0.746 | |
Chandra | 346 | 203 | 81 | 6.16 | 6.08 | 0.930 |
3090 | 281 | 56 | 7.82 | 8.11 | 0.956 | |
4369 | 318 | 63 | 8.66 | 8.97 | 0.950 | |
11586 | 106 | 42 | 2.86 | 2.88 | 0.884 | |
12039 | 197 | 115 | 5.11 | 5.17 | 0.936 | |
12040 | 150 | 86 | 3.63 | 3.68 | 0.911 | |
15580 | 390 | 134 | 2.31 | 2.37 | 0.932 | |
14222 | 327 | 103 | 10.6 | 9.47 | 0.953 | |
14223 | 115 | 30 | 8.61 | 9.02 | 0.948 | |
16478 | 183 | 71 | 5.71 | 4.88 | 0.965 | |
17574 | 128 | 63 | 4.19 | 3.55 | 0.942 |
We extracted the number of photon counts in our source region using the CIAO tool dmextract and the total exposure for the region using the same method on the exposure maps generated by NuSTARDAS nuexpomap for our NuSTAR data and CIAO tool mkexpmap for our Chandra data. To estimate (2) for Chandra, we similarly extract the counts and exposures of a corresponding background region and weight the background counts by the ratio of the background exposure to the source exposure. Unlike with Chandra, NuSTAR’s background is spatially asymmetric across detectors, so we could not use a background region separate from our source region. Instead, we set the background exposure to be the same as our source exposure, and used nuskybgd produced background maps to extract the background counts within our source region.
In general for both NuSTAR and Chandra the PSF (and therefore, the ECF) changes depending on the distance from the source to the optical axis. We use our extracted regions to measure the mean ECF for each observation. For Chandra, we generated simulated PSFs using the CIAO tool simulate_psf and MARX v5.4.0 and calculated the percentage of photons enclosed within the source region. For NuSTAR, we used NuSTARDAS product files to map the path of the optical axis for each observation and took the median x- and y-coordinate of its path. We then simulated the NuSTAR PSF based on the distance between the source and the optical axis and used this to calculate the ECF as was done for the Chandra observations.
3 Results
We present the fit results for our analyses in Tables 5 (NuSTAR epochs 1, 2, and all epochs combined) and 6 (Chandra all epochs combined and NuSTAR+Chandra combined fit). The results for the individual Chandra epochs are given in Appendix Table A1. In these tables we give the best-fit values and posterior distribution of the model parameters detailed in Table 2 for each data set as calculated through our BXA analysis. We report only on the most likely models for a given epoch or combined fit as determined by the Bayesian evidence (see below). In all but one case, the powerlaw, broken powerlaw, and double broken powerlaw were the three preferred models according to this metric. We report energy parameters in units of keV and normalizations in units of nJy. All other parameters are dimensionless.
Inst.-Epoch | Model | (nJy) | param. 1 | param. 2 | param. 3 (keV) | param. 5 | log(Z) | p(M—D) | /dof | |
NS-1 | PL | - | - | - | -0.12 | 0.271 | 37.8/49 = 0.771 | |||
Bkn. PL | - | -0.07 | 0.304 | 37.0/47 = 0.796 | ||||||
Dbl. Bkn. PL* | 0.00 | 0.357 | 36.7/45 = 0.823 | |||||||
NS-2 | PL* | - | - | - | 0.00 | 0.338 | 142.9/159 = 0.899 | |||
Bkn. PL | - | -0.03 | 0.315 | 141.6/157 = 0.902 | ||||||
Dbl. Bkn. PL | -0.01 | 0.330 | 141.7/155 = 0.914 | |||||||
NS-All | PL* | - | - | - | 0.00 | 0.373 | 196.9/211 = 0.933 | |||
Bkn. PL | - | -0.08 | 0.310 | 196.0/209 = 0.938 | ||||||
Dbl. Bkn. PL | -0.08 | 0.310 | 201.5/207 = 0.974 |
We compared the alternative models based on three different metrics which are suitable for comparing non-nested models. The first is the Bayesian evidence, defined as the integral over the parameter space: , where is the parameter vector. The second is the Bayesian information criterion, (Schwarz, 1978), where is the log-likelihood value, is the number of data points, and is the number of degrees of freedom. The BIC is an approximation under the assumption that the posterior distribution is strongly single-peaked, which is well-supported by the posterior distributions for our data. The final comparison statistic we use is the Akaike information criterion, (Akaike, 1974). The is a more conservative metric which measures the information loss of a specific model and is useful for determining if a model has the appropriate number of degrees of freedom. In all cases included in this study, comparing models using and only reinforces the paradigm established by the Bayesian evidence. As such, we omit these values from Tables 5 and 6. Although we emphasize that it is not useful for comparison nor is it used in fitting at any point, we do calculate the for each model as a check of the goodness-of-fit. This value is calculated by loading the best-fit model from the BXA analysis into XSPEC and binning the spectra to at least 30 counts per bin using ftgrouppha. The Bayesian evidence, the relative probabilities of the different models (which are calculated directly from the Bayesian evidence), and the are all recorded in Tables 5 and 6.
Inst.-Epoch | Model | (nJy) | param. 1 | param. 2 | param. 3 (keV) | param. 5 | log(Z) | p(M—D) | /dof | |
CH-All | PL | - | - | - | -0.78 | 0.082 | 464.0/598 = 0.776 | |||
Bkn. PL | - | -0.08 | 0.413 | 462.6/596 = 0.776 | ||||||
Dbl. Bkn. PL* | 0.00 | 0.496 | 453.9/594 = 0.764 | |||||||
CH+NU-All | PL | - | - | - | -1.41 | 0.024 | 660.9/802 = 0.824 | |||
Bkn. PL | - | 0.00 | 0.629 | 462.6/596 = 0.776 | ||||||
Dbl. Bkn. PL | -0.26 | 0.346 | 453.9/594 = 0.764 |
3.1 NuSTAR spectral fitting
According to the Bayesian evidence resulting from our analysis of the combined NuSTAR data (recorded in Table 5), the simple power-law is the slightly preferred model for the 3-20 keV band. (The broken power law and double broken power law appear at first acceptable if slightly less-preferred models, but see discussion of parameter values below.) The exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence. The simple power law model is also favored when comparing the and for the different models.
The best-fit folded power law model and posterior range are plotted alongside the NuSTAR FPMA and FPMB data in the top panel of Figure 5, and the residuals of the model are plotted in the bottom panel. This simple power law fit of the 3-20 keV spectrum has a best-fit photon index of and a normalization of nJy at 1 keV. The maximum likelihood value for the constant offset factor we introduce to account for sensitivity differences between the FPMA and FPMB modules is which is in agreement with previous estimates of a calibration difference between the two detectors (Madsen et al., 2015).
We find no evidence of a spectral break or cutoff in our NuSTAR X-ray data alone. The maximum likelihood parameter values for the broken powerlaw and double broken powerlaw models are both asymptotic approximations of the simple power law. That is, the posterior distribution of the broken power law is strongly peaked at a parameter distribution such that and the posterior distribution of the double broken powerlaw is strongly peaked such that .
We also fit the 2 epochs of NuSTAR data separately. The Bayesian evidence of the second epoch (which is 3 times the exposure of the first) again favored the simple powerlaw model, with the best-fit broken powerlaw and double broken powerlaw models being slightly less probable and showing little change in spectral index across the energy band. For the first epoch, the double broken powerlaw is actually the slightly preferred model, however the parameter best-fit values and errors make it clear that it is effectively equivalent to a single flat power-law spectrum, with , , and .
Assuming a simple powerlaw as the correct model, the photon index showed no significant variation between the two major epochs, ranging from for epoch 1 to for epoch 2. However, the best-fit normalization changed substantially between the two epochs, with an approximately increase in value from epoch 1 to epoch 2, suggesting variability in the hard X-ray flux. This is further examined and quantified in Section 3.5.
3.2 Chandra X-ray spectrum
According to the Bayesian evidence resulting from our analysis of the combined Chandra data (recorded in Table 6), the double broken power-law is the slightly preferred model for the 0.5-7 keV band (although see discussion below). The broken power law and simple power law are also acceptable if slightly less-preferred models, while the exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence.
Interestingly, when we fit the individual epochs of Chandra data separately, the simple powerlaw (1, 3-5) and broken powerlaw (2, 6, 7) are consistently favored over the double broken powerlaw (albeit only slightly). The preference for the double broken powerlaw in the combined fit might be the result of variations from epoch to epoch which manifest as an additional spectral break in the combined data set. Because of this, we instead opt to consider the broken power law model as the “true” spectral model as it is only slightly less preferred than the double broken powerlaw in the combined dataset.
The best-fit broken power law model has a primary photon index of , a secondary photon index of , a break energy at , and a normalization of nJy at 1 keV. This folded model and its posterior range are plotted alongside the combined 0.5-7 keV Chandra data in the top panel of Figure 6, and the residuals of the model are plotted in the bottom panel. These results are in rough agreement with Hardcastle et al. (2016), whose analysis of the Chandra observations of the Western hot spot also slightly preferred the broken power law model fit to the spectrum.
Similar to our NuSTAR analysis, comparing the changes in the best-fit single powerlaw models between various Chandra epochs reveals that the photon index varies little, ranging from to . Also similarly, albeit smaller in scale, we see a roughly 10 decrease in the normalization from Chandra epoch 4 to epoch 5, or a 2.33 change. This is similar to the variability previously reported by Hardcastle et al. (2016). The values of the normalization and integrated flux in the 3-8 keV flux band shared by both NuSTAR and Chandra across the different epochs of data from both telescopes are plotted in Figure 7, with the photon index in the lower panel. Further quantification and discussion of the X-ray variability is continued in Section 3.5.
3.3 Combined Chandra + NuSTAR X-ray spectrum
According to the Bayesian evidence resulting from our analysis of the combined fit of our Chandra+NuSTAR data (recorded in Table 6), the broken power-law is the preferred model for the 0.5-20 keV wide band. The double broken power law and simple power law are also acceptable if less-preferred models, while the exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence.
The best-fit broken power law model and posterior range are plotted alongside the combined Chandra and NuSTAR data in the top panel of Figure 8, and the residuals of the model are plotted in the bottom panel. This broken power law model has a best-fit photon indices of and , a break energy of , and a normalization of nJy at 1 keV.
Although the Bayesian evidence suggests that the broken powerlaw and double broken powerlaw models are significantly more probable than the simple powerlaw model, a closer examination of the results suggest they must be interpreted with caution. Firstly, the best fit primary break energy value for both the broken and double broken powerlaw models is approximately keV, close to the lower boundary of the spectral range and differing from the value of 3 keV for the Chandra fit alone. The photon index after the break is approximately , which is in agreement with the best fit simple powerlaw photon index calculated for most epochs of both NuSTAR and Chandra data. In the case of the double broken powerlaw model, the second spectral break shows no significant change in spectral index for before and after the second break ( to ), and thus can be disregarded. Importantly, the joint fit also includes a relative normalization factor for NuSTAR to force agreement with Chandra; the value is , a larger difference than expected for instrumental effects alone.
3.4 Spectral energy distribution
For comparison and to produce the most complete spectrum of Pictor A’s western hot spot to date, we collected multifrequency archival data from previous studies, including radio and infrared observational data from the VLA (Meisenheimer et al., 1997), observations from Spitzer (Werner et al., 2012), WISE (Isobe et al., 2017), and SPIRE (Isobe et al., 2020), as well as additional archival data recorded in the aforementioned studies.
The full spectral energy distribution (SED) of the western hot spot is shown in Figure 9. We include the model of Isobe et al. (2020) to describe the low energy emission, consisting of a primary cutoff powerlaw and an ‘excess’ broken cutoff powerlaw to account for the observed flux at mid- and far-infrared frequencies. We additionally plot as calculated from the unfolded data of our independent Chandra and NuSTAR datasets. The data points are corrected for background, instrument response, and absorption using XSPEC’s plot ufspec command.
There is an offset in normalization between our NuSTAR and Chandra data which is evident both in the SED and in the best-fit parameter values recorded in Table 6, suggesting that the hotspot may have been in a somewhat higher state during the recent NuSTAR observation. Because the XSPEC model by default normalizes the flux at 1 keV (technically below the NuSTAR sensitivity range), this apparent variability likely explains the apparent shift of the spectral break energy in the combined fit as an artifact of the forced agreement. A visual inspection of the SED suggests that there is indeed a slight break at around 2 keV. At higher energies, the spectral fits for essentially all data sets here considered agree on a very nearly perfectly flat spectrum with no other break or cutoff.
3.5 Variability analysis
We find strong evidence of statistically significant variability in the total Chandra-measured X-ray flux for the western hot spot. When we include all 11 Chandra observations, the returned p-value is , allowing us to reject the steady rate hypothesis at a significance of , approximately twice the significance over previous reports. We also analyzed smaller subsets of observations, in order to constrain the shortest timescale of the flux variability. For example, applying our variability analysis to only epochs 5 and 6 of the Chandra data, which are separated in time by approximately 2 months, still gives a p-value of , or . This suggests significant X-ray flux variations on the timescales of months, which by light-crossing time arguments would place an upper limit of pc on the varying region. This result also supports the findings of Hardcastle et al. (2016), who reported a change in the flux of the western hot spot of approximately on a timescale of approximately 1 month.
For the variability analysis of the NuSTAR data alone, we recorded the necessary data outlined in Section 2.6 separately for both the FPMA and FPMB detectors. We then generated a combined data set, using the total photon counts, total exposure, and average ECF from the two detectors, which are presented in Table 3. We then analyzed this combined data set using our maximum likelihood analysis, which returns , equivalent to .
The two epochs of NuSTAR data alone do not demonstrate highly significant variability, but when we include the full 22 year time-span of NuSTAR and Chandra in the shared 3-8 keV energy band (data recorded in 4), our maximum likelihood method returned a p-value below the machine precision value of , , giving a significance of at least . This result for the Western hotspot is the strongest evidence yet of short-timescale (years) variability in the Pictor A jet. These results, along with additional tests for short-term variability are summarized in Table 7, where we list the instrument(s), epochs, resultsing p-values and equivalent significance, and the longest time baseline of the data set.
Instrument | Epochs | p-value | Significance () | (t) |
Chandra | 1-3 | 0.064 | 1.85 | 9 yr |
Chandra | 5,6 | 9.02e-4 | 3.32 | 3 mo |
Chandra | 4-7 | 3.19e-5 | 4.16 | 2 yr |
Chandra | All | 2.04e-11 | 6.70 | 15 yr |
NuSTAR | 2 | 0.820 | 0.23 | 1 wk |
NuSTAR | All | 0.082 | 1.74 | 6 yr |
NuSTAR+Chandra | All | 2.2e-16 | 8.21 | 22 yr |
4 Discussion
The origins of the X-ray emission in AGN hot-spots has been debated extensively in the literature (e.g., Tavecchio et al., 2000; Celotti et al., 2001; Hardcastle et al., 2004; Georganopoulos et al., 2006; Tingay et al., 2008; Migliori et al., 2020). Initial studies of Pictor A in particular suggested that a standard one-zone SSC model was incapable of explaining the broadband SED, including an unusually high X-ray-to-radio flux ratio (Wilson et al., 2001). As a result, numerous alternatives were considered, most notably the IC-CMB model (Tavecchio et al., 2000; Celotti et al., 2001).
The shape of the X-ray spectrum can potentially rule out IC/CMB under certain assumptions. The IC model requires the upscattered emission to follow the spectral shape of the primary, radio-emitting synchrotron curve shifted over in frequency and luminosity (Georganopoulos et al., 2006). Therefore, we can test the IC-CMB hypothesis by comparing the shapes of the spectral curves for the different emission components. In particular, if IC-CMB is responsible for the anomalous X-ray emission we expect to find that the X-ray spectral index is roughly equivalent to the low-frequency (MHz) radio spectral index. However, we are generally unable to measure the spectral index at these frequencies due to a lack of observations. We must instead compare the X-ray spectral index and the GHz radio-optical spectral index, which we naively assume is roughly equivalent to or softer than the MHz radio spectral index. We thus expect to find that the X-ray spectral index is equal to or harder than this GHz radio-optical spectral index. In contrast, if the anomalous X-rays are produced by synchrotron radiation, the X-ray spectrum is not subject to the same requirements, and we might observe an X-ray spectral index which is softer than the GHz radio-optical spectral index (Georganopoulos et al., 2006; Jester et al., 2006).
The results of this analysis (in agreement with previous, e.g. Hardcastle et al., 2016; Thimmappa et al., 2020) strongly disfavor an IC-CMB origin for the X-ray photons, based on the variability as well as the mismatch between the X-ray spectral index () and the radio spectral index (). This variability timescale disfavors an IC-CMB origin for the X-ray emission, as IC-CMB predicts cooling timescales of the order of yrs (Georganopoulos et al., 2006). In contrast, the lifetime of an electron emitting synchrotron radiation of characteristic energy , gyrating in a magnetic field of strength is given by . Under the hypothesis of localized substructures within the hotspot of less than a kpc in radius, we expect a magnetic field on the order of (Isobe et al., 2017). These values result in a cooling timescale which is consistent with the observed variations in X-ray flux.
The shortest timescales of variability in the Chandra data band suggest an emitting region of less than a pc in radius. These findings support the proposal of Tingay et al. (2008), who studied the Pictor A western hot spot using the Very Long Baseline Array (VLBA) and resolved pc-scale components within the hot spot. They suggested the X-rays could be produced by discrete substructures within the hot spot which are highly magnetized, which also could explain the observed excess emission in the IR (Isobe et al., 2017, 2020). The lack of high spatial resolution in the X-rays prevents us from spatially resolving the variability, but clearly the total flux we observe is integrated over many discrete regions, and therefore the low level (tens of percent) variability we see in aggregate translates to much larger-amplitude flaring within the actual small-scale region undergoing variability. Future coordinated monitoring campaigns with e.g., the VLBA and a next-generation sensitive and high-spatial-resolution X-ray telescope such as AXIS may reveal these sites and the particle acceleration within them in action.
5 Conclusions
We have presented the results of a comprehensive analysis of the Western hot spot of Pictor A using a recent deep NuSTAR observation combined with archival data from both NuSTAR and Chandra. We presented seperate (soft/hard) and combined fits to the wide X-ray band (0.5-20 keV) an danalyzed the X-ray emission for variability using the maximum likelihood method of Meyer et al. (2023). The main results of these analyses are:
(i) We found that the NuSTAR data alone is best fit by a simple power law of photon index and a normalization at 1 keV of . We do not find any evidence for a spectral break or cutoff energy in the hard X-ray band. The difference in normalization between the recent 2021 observations and previous epochs of both Chandra and NuSTAR data is suggests variability within the western hot spot.
(ii) We found that the Chandra data alone is best fit by a broken power law of photon indices and , a break energy at , and a normalization of nJy at 1 keV.
(iii) We found that the combined Chandra and NuSTAR data are best fit by a broken power law model, with a best-fit photon indices of and , a break energy of , and a normalization of nJy at 1 keV. However, the fit requires a significant offset factor of for NuSTAR relative to Chandra. This suggests that the hot spot flux may have been higher during the recent NuSTAR epoch and that interpretations of the joint fit must be made cautiously.
(iv) The maximum likelihood variability analysis of the Chandra observations of the western hot spot finds statistically significant () variations in the 0.4-8 keV X-ray emission, confirming the results of Hardcastle et al. (2016) and Thimmappa et al. (2020). We also confirm a significant () variation on a timescale of months, placing an upper limit on the size of the varying region of pc.
(v) The variability analysis of the NuSTAR data alone is unable to find evidence of statistically significant variability in the hard, 3-20 keV X-ray flux, with a maximum significance of . However, analyzing both the NuSTAR and Chandra data in the shared 3-8 keV energy range gives the most significant detection of variability yet, at .
Collectively, these results suggest a synchrotron origin for the hot spot X-ray emission, as the alternative beamed IC-CMB model is incapable of explaining either the rapid X-ray variability or the discrepancy between the measured X-ray spectral index () and the expected low-frequency radio spectral index ().
The emerging picture of the hot spot is of multiple localized, highly magnetized, pc-scale substructures emitting X-ray predominantly via synchrotron radiation. This paradigm is capable of explaining the rapid X-ray variability and the difference in spectral index between the radio and X-rays. Magnetic reconnection models are capable of reproducing the necessary in situ acceleration in smaller regions.
Follow-up observations of Pictor A with NuSTAR might allow for a more significant detection of variability in the hard X-ray band and could confirm the flaring detected in the recent observations. The next major evolution will be the study of other nearby hot spots in the hard X-ray band. As mentioned previously, this is not a feasible task for NuSTAR due to its angular resolution and the lower X-ray flux of most other hot spots relative to Pictor A’s western hot spot. Next generation X-ray missions with higher angular resolution and sensitivity such as FORCE (Nakazawa et al., 2018), HXMT (Zhang et al., 2020), and Athena (Barret et al., 2023) should be capable of analyzing whether Pictor A’s western hot spot is an outlier or a typical representative of the X-ray jet population.
Appendix A NuSTAR background models
We briefly detail the various background components here; a more complete overview can be found in (Wik et al., 2014).
Internal background. This element models radiation resulting from the environment of NuSTAR’s orbit and consists of two major components. The first is a featureless continuum, modeled as a broken power law with a break at 124 keV which dominates the background above 30 keV. The second is a collection of fluorescense and activation lines which dominates from approximately 20-30 keV, with weaker lines present up to keV.
Aperture background. The NuSTAR optical and focal plane modules are separated by an unenclosed mast. As a result, far away, off-axis stray light can leak into the aperture stop and strike the detector. Some of this stray light is blocked by the optics bench, resulting in a variable, spatially non-uniform background gradient across the detectors. The cosmic X-ray background (CXB) is the main contributor to this ’aperture’ background, which dominates the background spectrum below 20 keV. nuskybgd uses the canonical HEAO-1 A2 spectral model for the CXB (Gruber et al., 1999; Revnivtsev et al., 2003; Churazov et al., 2007).
Scattered background. The NuSTAR detectors are also affected by reflected and scattered X-rays from parts of the observatory. The main sources of these X-rays are the CXB, the Sun, and the Earth. These radiation sources are subdominant at all energies, but can contribute significantly at low energies. In particular, during periods of heightened solar activity, the solar background appears as a 1 keV thermal spectrum which can rival the aperture background radiation flux in the 3-5 keV energy range.
fCXB. The ‘focused’ CXB (fCXB) is a background radiation component resulting from unresolved sources within the FOV. The contribution of the fCXB is subdominant at all energies, but can contribute significantly below 15 keV, with a magnitude of approximately 10 of the CXB flux.
The nuskybgd package uses XSPEC333https://heasarc.gsfc.nasa.gov/xanadu/xspec (Arnaud, 1996) to fit the spectrum of the entire background across the NuSTAR detector modules, taking into account each component and any spatial variations present. nuskybgd groups the input spectrum to a minimum of 30 counts per bin and calculates the parameter normalizations using the statistic. Within XSPEC, the background model is defined using a combination of four different models. apbgd and fxapbgd model the ‘aperture’ and ‘focused aperture’ or ‘fCXB’ components and are both defined as cutoff power-law components (cutoffpl) with a photon index of and a high-energy cutoff of keV. The line_bgd component models the fluorescense and activation line elements of the ‘internal’ background component and the Solar thermal emission of the ‘scattered’ background. It is comprised of 29 different Lorentzian (lorentz) line profiles and a thermal plasma component (apec) with temperature . The final element is the particle_bgd component which models the featureless continuum of the ‘internal’ background as a broken power-law with photon indices of and a break enery of keV.
We defined the background using 3 concentric annular regions which excluded emission from Pictor A and its jet and hot spot. An image example of these annular background regions for one of our NuSTAR observations is shown in Figure A1. The spectral parameters of each model component such as the power-law photon indices and the line energies and widths are set by nuskybgd and fixed prior to fitting and do not change from observation to observation. Only the normalizations of these components are free to vary during the background fitting process. The normalizations of each component are recorded for each observation in Supplemental Table 1 in units of ph.
Appendix B Results for individual Chandra epochs
In the main manuscript we have reported the results of our analysis of the western hot spot spectrum using NuSTAR data alone in Table 5 and the analysis of the total Chandra and the combined NuSTAR+Chandra datsets in Table 6. Due to the number and length, the results for individual Chandra epochs are reported here in Appendix Table A1.
Inst.-Epoch | Model | (nJy) | param. 1 | param. 2 | param. 3 (keV) | param. 5 | log(Z) | p(M—D) | /dof | |
CH-1 | PL* | - | - | - | 0.00 | 0.416 | 62.4/69 = 0.904 | |||
Bkn. PL | - | -0.05 | 0.370 | 61.1/67 = 0.912 | ||||||
Dbl. Bkn. PL | -0.34 | 0.190 | 63.6/65 = 0.979 | |||||||
CH-2 | PL | - | - | - | -0.50 | 0.156 | 116.1/160 = 0.726 | |||
Bkn. PL* | - | 0.00 | 0.492 | 112.1/158 = 0.709 | ||||||
Dbl. Bkn. PL | -0.19 | 0.318 | 110.6/156 = 0.709 | |||||||
CH-3 | PL* | - | - | - | 0.00 | 0.348 | 97.2/124 = 0.784 | |||
Bkn. PL | - | -0.03 | 0.325 | 97.7/122 = 0.801 | ||||||
Dbl. Bkn. PL | -0.09 | 0.283 | 97.4/120 = 0.812 | |||||||
CH-4 | PL* | - | - | - | 0.00 | 0.373 | 20.7/24 = 0.862 | |||
Bkn. PL | - | -0.02 | 0.356 | 19.9/22 = 0.904 | ||||||
Dbl. Bkn. PL | -0.18 | 0.246 | 21.0/20 = 1.051 | |||||||
CH-5 | PL* | - | - | - | 0.00 | 0.391 | 38.9/75 = 0.519 | |||
Bkn. PL | - | 0.00 | 0.365 | 39.6/73 = 0.543 | ||||||
Dbl. Bkn. PL | -0.24 | 0.225 | 39.5/71 = 0.556 | |||||||
CH-6 | PL | - | - | - | -0.08 | 0.304 | 43.1/63 = 0.684 | |||
Bkn. PL* | - | 0.00 | 0.365 | 42.7/61 = 0.700 | ||||||
Dbl. Bkn. PL | -0.35 | 0.163 | 41.9/59 = 0.711 | |||||||
CH-7 | Exp. Cutoff PL | - | - | -0.02 | 0.294 | 59.7/70 = 0.852 | ||||
Bkn. PL* | - | 0.00 | 0.308 | 57.9/69 = 0.839 | ||||||
Dbl. Bkn. PL | -0.24 | 0.177 | 57.2/67 = 0.854 |
Appendix C Image deconvolution with LIRA
To detect the inner jet or any features in the western hot spot, for example the “bar” seen in Chandra images, possibly buried in the raw NuSTAR observations, we deconvolved the merged sky image of Pictor A (see Figure 2) with an effective point spread function (PSF).
We use a statistical tool called Low Count Image Reconstruction and Analysis (LIRA, Esch et al., 2004; Stein et al., 2015) to deconvolve images. Although LIRA was originally designed to detect the statistical significance of structures in low-count high-energy observations, it also has been used to deconvolve sky images (e.g., Kashyap et al., 2017; Reddy et al., 2021b, 2023). We refer the reader to Stein et al. (2015) for a full statistical treatment. Briefly, LIRA models the brightness distribution of an observation using two model images. One is a baseline with known sources of emission (e.g., core and background), which is specified by the user, and an added model, which is to be inferred containing any emission over the baseline, required to explain the observation. We can reconstruct the observation by convolving their sum with the PSF. LIRA performs Bayesian analysis using a flexible multi-scale representation of the observation to infer the posterior distribution of the added model and samples it with a Markov Chain Monte Carlo (MCMC) sampler to generate a series of images. These images represent the observation’s brightness distribution after subtracting the baseline. With nuskybgd-generated background as the baseline, we generate 2000 images after discarding the first 1000 images as burn-in and average them to produce a deconvolved image.
Due to the spatial dependence of NuSTAR’s PSF, using a single PSF to deconvolve both the core and the hot spot, which are separated by 4’, may lead to artifacts in the output images. Hence, we separately deconvolve the core and the hot spot regions. Furthermore, to determine the energy dependence of the structure in the jet and the hot spot, we divide the image into two energy ranges of 3-12 keV and 12-78 keV, as shown in the top and bottom rows of figure A2. The left and right two panels in each row show core and WHS images, respectively. We set the core image size to 128x128 and WHS to 64x64, following the input requirements of LIRA.
The top and bottom rows of panels in figure A2 show the sky and LIRA-deconvolved images of the core and the WHS in the energy ranges 3-12 keV and 12-78 keV, respectively. We find no evidence for an inner jet in the core’s deconvolution in both the energy ranges (panels b and f), indicating its emission is within the error limits of the core’s PSF and background. This absence of emission is expected in the sense that the flux from the inner jet is close to the detection limit of NuSTAR. The 0.5-7 keV energy flux of the brightest feature in the inner jet is erg cm s, while the 3-sensitivity of NuSTAR with a 1 Ms exposure444Measured in the 6-10 keV band (about three times larger than our combined 300 ks exposure) is erg cm s (Harrison et al., 2013). Finally, because we also find no evidence for a jet beyond 12 keV (fig. A2(e)), we use the estimated 10-30 keV sensitivity of erg cm s as an upper limit for each knot. Hence, we take erg cm s as an upper limit on the inner jet’s 10-30 keV flux.
Panels (d) and (f) show the LIRA-deconvolved images of the WHS in the 3-12 keV and 12-78 keV bands, respectively, where the detected structure appears spatially correlated with previously known features. The 3-12 keV images show point sources at the southern end of the bar and the hotspot, consistent with the structure found in Chandra images. Interestingly, the 12-78 keV band lacks a point source near the hotspot, but we find a bright point feature spatially coincident with the bar’s southern peak. This feature indicates the jet presumably produces an electron population capable of generating hard X-rays when it enters the turbulent hotspot region. However, the limited statistics in the bar region preclude a detailed spectral analysis and require a much deeper observation to understand the true nature of this hard X-ray emission.